OpenStudy (anonymous):

find the derivative of f(x)= |x-3| +2 help plsss :D

OpenStudy (anonymous):

depends on the domain, since the absolute value is a piecewise function

OpenStudy (anonymous):

if \(x>3\) then \(|x-3|=x-3\)

OpenStudy (anonymous):

whereas if \(x<3\) then \(|x-3|=3-x\)

OpenStudy (anonymous):

so your derivative is either 1 or -1 depending on whether \(x>3\) or \(x<3\)

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle \frac{d}{du}\left[~|u|~\right]=\frac{u}{|u|} \cdot u' }\)

OpenStudy (solomonzelman):

you can prove that, by using the definition \(\large\color{black}{ \displaystyle |u|=\sqrt{u^2}}\)

OpenStudy (anonymous):

alright that makes sense, do you guys know the mean value theorem

OpenStudy (anonymous):

heard of it yes

OpenStudy (anonymous):

im given the interval [1,6] and i need to figure out which hypothesis is not met

OpenStudy (anonymous):

is it that f'(x) is not differentiable?

OpenStudy (anonymous):

ok that is a reasonable question we have to check that the function is continuous on the closed interval and differentiable on the open one

OpenStudy (anonymous):

don't be confused by that differentiable doesn't mean "i can find the derivative" differentiable means "the derivative exists on the entire interval"

OpenStudy (anonymous):

what is the function?

OpenStudy (dan815):

Given a function F(x) in some interval [a,b], a<=c<=b there exists a value c, such that F(c)= (F(b)+F(a))/2

OpenStudy (anonymous):

yea but it doesnt right? cuz we just said its either 1 or -1?

OpenStudy (anonymous):

oooh it was the function above!

OpenStudy (anonymous):

yes, the answer is NO because your function is not differentiable at \(x=3\)

OpenStudy (anonymous):

alright thanks for clearing things up :D

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

@dan815 over two?

OpenStudy (dan815):

yeah its the average value

OpenStudy (anonymous):

oh no that is not what the mvt says at all

OpenStudy (anonymous):

isnt it f(b)-f(a)/b-a = f'(c) for the mean value theorem?

OpenStudy (anonymous):

in fact, not to be argumentative, but it is not even close

OpenStudy (anonymous):

@Noodles24 that is the CONCLUSION of the mvt

OpenStudy (dan815):

oh okay srym b-a i see

OpenStudy (anonymous):

yea i thought thats what he was trying to say

OpenStudy (anonymous):

if \(f\) is continuous on \([a,b]\) and differentiable on \((a,b)\) then there is a \(c\in (a,b)\) with \[f'(c)=\frac{f(b)-f(a)}{b-a}\]

OpenStudy (dan815):

i misunderstood what MVT was for again hehe, So i guess this means we are saying there is a F(c) such that it is the average change of the function on the interval [a,b] I previously thought MVT is just the average value not the average change xD

OpenStudy (anonymous):

i got another quick question for ya

OpenStudy (anonymous):

if the derivation of something is just like one number, such as 1, is it differentiable?

OpenStudy (anonymous):

idk if u would ever have that though

OpenStudy (dan815):

yeah if its non differentiable, we can have like an infinity or some un accounted change of the average value showing up

OpenStudy (anonymous):

alright i got another kinda stupid question

OpenStudy (anonymous):

cos(x)-xsin(x)=0

OpenStudy (anonymous):

how u solve for x lol, i feel like this is something easy but i cant seem to rememeber

OpenStudy (dan815):

x=cos(x)/sin(x) 1/x=tan(x) so 1 over the angle = opp/adj

OpenStudy (dan815):

or adj/opp = x

OpenStudy (anonymous):

i understand but i need like a numerical value for x

OpenStudy (dan815):



OpenStudy (dan815):

yeah can u see wheere this would ben true on a unit circle

OpenStudy (dan815):

ah im lagging cant really see my text

OpenStudy (anonymous):

alright i think i understand it, thanks for all the help man <3

OpenStudy (dan815):

ye sure :)